OpenStudy (anonymous):

what is the product of 5 sqrt 25d* 2d sqrt3? simplify if possible. 5dsqrt3 10dsqrt5d 25dsqry6d 50dsqrt3d

OpenStudy (anonymous):

help me please i only have half hour

OpenStudy (anonymous):

is it \[\sqrt{25 d} \]

OpenStudy (anonymous):

or \[\sqrt{25}\]d

OpenStudy (anonymous):

@rajfk35 Am I understanding what you meant here? Can you follow my steps? \[5 \sqrt{25}d* 2d \sqrt{3} = 5 * 2 * d * d * \sqrt{25} * \sqrt{3} = ?\]

OpenStudy (anonymous):

\(\sqrt{25}=\sqrt{5*5}=5\) 5 * 2 * 5 = 50 d * d = d\(^2\) and then \(\sqrt3\) See how that works?

OpenStudy (anonymous):

Or did you mean: \[5 \sqrt{25d}* 2d \sqrt{3} = 5 * \sqrt{d} * \sqrt{25} * 2* d * \sqrt{3} = ?\]

OpenStudy (anonymous):